This has to do with a quadratic function.
For the quadratic function f(x)=5.92x^2-3.135x-5.86 , find the maximum or minimum value, x- intercepts and the y - intercept.
Does the quadratic have a maximum or minimum?

Answer by venugopalramana(3286)

This has to do with a quadratic function.
For the quadratic function f(x)=5.92x^2-3.135x-5.86 ,
F(X)=5.92=5.92(X-0.26478)^2-6.27504
HENCE THE FUNCTION HAS MINIMUM AT X=0.26478....AND THE MINIMUM VALUE IS
F(X)=-6.27504
X INTERCEPT IS AT F(X)=0
X=0.26478+ OR - SQRT(6.27504/5.92) = 1.29433 OR -0.76477
Y INTERCEPT IS AT X=0
F(X)=5.92*0.26478^2-6.27504= -5.86

find the maximum or minimum value, x- intercepts and the y - intercept.
Does the quadratic have a maximum or minimum?